OpenStudy (anonymous):

Please Help Me!!! Will Medal!!! At a local restaurant, a new dinner menu is available that offers 4 courses - an appetizer, a salad, a main dish, and a dessert - all for one price. There are 5 appetizers, 2 salads, 3 main dishes, and 5 desserts. How many dinner outcomes are possible? A. 15 B. 150 C. 5 D. 24

OpenStudy (kropot72):

Each of the 5 appetizers can be combined with each of the 2 salads, giving 5 * 2 = 10 choices. Each of these 10 choices can be combined with each of the 3 main dishes, giving 10 * 3 = 30 choices. Each of these 30 choices can be taken with each of the 5 desserts giving a total of 30 * 5 = 150 possible outcomes.